OpenStudy (anonymous):

HELP ME FOR THE BEST TESTIMONY OF YOUR LIFE: The area of the region bounded by the curves y = f(x) = 2x^3 – 6x^2 – 2x + 6 and y = g(x) = –x^3 + 3x^2 + x – 3 is given by: ?

OpenStudy (anonymous):

Refer to the Mathematica attachement which includes a plot.

OpenStudy (anonymous):

Sorry, there were some misspellings in the attachment.

OpenStudy (anonymous):

What is your answer?

OpenStudy (anonymous):

Is this correct?

OpenStudy (anonymous):

\[\left\{\int\limits_{-1}^1 \left(3 x^3-9 x^2-3 x+9\right) \, dx,\int\limits_1^3 \left(-3 x^3+9 x^2+3 x-9\right) \, dx\right\}\]Both integrals are equal to 12. \[\{12,12\}\]

OpenStudy (anonymous):

Mathematica likes to leave terms in reverse order.

OpenStudy (anonymous):

Can you do this one for me? Find the area between the curves y = f(x) = x^3 – 3x and y= g(x) = x (the square root of 4-x^2) on the interval 0 ≤ x ≤ 2.

OpenStudy (anonymous):

Find the area between the curves y = f(x) = x^3 – 3x and y= g(x) = x (the square root of 4-x^2) on the interval 0 ≤ x ≤ 2. a)8/3 b)20/3 c)4/3 d)44/3 e)14/3 Which is it? :/

OpenStudy (anonymous):

The second solution is attached. I worked the problem to get answer e) The curves cross at about x = 1.9 and the area goes negative after that x value.